OpenStudy (anonymous):

Find the value of y so that the line passing through (6, y) and (7, 5) has a slope of -3

OpenStudy (tommo_lcfc):

I have just attempted this question successfully. Just use the equation for a straight line (not y = mx + c, the other one!) and find the equation using the known point. Then use the other point to find the value of y. I can take you through it if needed. :)

OpenStudy (anonymous):

please do

OpenStudy (anonymous):

take me through it

OpenStudy (tommo_lcfc):

Gladly! :) We know the gradient of the straight line concerned is -3. The line passes through the point (7,5). Use y - y1 = m(x - x1) with y1 as 5 and x1 as 7 to get the straight line equation. When you have that, substitute the other value of x (i.e. 6) into this equation to get the unknown y-value. The value of y should be 8, unless I have made a mistake somewhere. :)

OpenStudy (anonymous):

help mee